1/3x-2/5=1/4(x+3) solve for x
Answer by rfer(16322)

1/3x-2/5=1/4x+3/4
4/12x-3/12x=15/20+8/20
1/12x=23/20
x=(23/20)/(1/12)
x=23/20*12/1
x=276/20
x=13 4/5